OpenStudy (anonymous):

Can someone explain why taking the derivative of 3(e^x + e^-x) = 3(e^x - e^-x)

OpenStudy (anonymous):

f(x) = \[3(e^x +e ^{-x}) \] f'(x) = \[3(e^x - e ^{-x}) \]

OpenStudy (solomonzelman):

I would expand and derive term by term.

OpenStudy (solomonzelman):

where you have e^(-x), apply the chain rule for the exponent, but whatever you have in the chain for the exponent, you multiply it on the outside, not inside the exponent.

OpenStudy (solomonzelman):

I mean differentiate term by term, not derive. (I get critised when I use word derive in this sense)

OpenStudy (solomonzelman):

\(\large\color{blue}{f(x)=3e^{x}-3e^{-x}}\)

OpenStudy (fibonaccichick666):

so what is the derivative of e^{-x}?

OpenStudy (solomonzelman):

\(\large\color{blue}{f'(x)=\frac{d}{dx}3e^{x}-\frac{d}{dx}3e^{-x}}\) \(\large\color{blue}{f'(x)=3\frac{d}{dx}e^{x}-3\frac{d}{dx}e^{-x}}\) \(\large\color{blue}{f'(x)=3e^{x}-3\frac{d}{dx}e^{-x}}\)

OpenStudy (solomonzelman):

e^x remains e^x.

OpenStudy (solomonzelman):

\(\large\color{blue}{\frac{d}{dx}e^{-x}=e^{\color{red}{-x}}\times\frac{d}{dx}(\color{red}{-x}) }\)

OpenStudy (anonymous):

Ok so that would be \[-1 e ^{-x}\]

OpenStudy (fibonaccichick666):

yes,

OpenStudy (solomonzelman):

yes, which is \-e^(-x)

OpenStudy (anonymous):

Ok so thats why...

OpenStudy (anonymous):

Thank you!!!

OpenStudy (solomonzelman):

So we get, \(\large\color{blue}{ f'(x)=3e^x-(-3x^{-x}) }\) \(\large\color{blue}{ f'(x)=3e^x+3x^{-x} }\)

OpenStudy (anonymous):

Thanks I get it now :))))

OpenStudy (solomonzelman):

So when the initial function is, \(\large\color{blue}{ f(x)=3e^x-3x^{-x} }\) then \(\large\color{red}{ f'(x)=3e^x+3x^{-x} }\) and if, the initial function is, \(\large\color{blue}{ f(x)=3e^x+3x^{-x} }\) then \(\large\color{red}{ f'(x)=3e^x-3x^{-x} }\)

OpenStudy (solomonzelman):

right, just the chain rule, derivative of -x, you multiply times -1.

OpenStudy (anonymous):

Yep, learning this stuff one step at a time haha.Thanks! )))

OpenStudy (solomonzelman):

anytime:) I was once new to this too:) But it gets easy to differentiate any function, unlike integration.